# SAM-Assisted Remote Sensing Imagery Semantic Segmentation with Object and Boundary Constraints

### 05 Dec 2023 v1

@@ III-B1 Object Loss
where $\mathcal{G}$ calculates the sum of all pixels in the spatial dimension and reshapes to its original shape while $N^{i}$ is the number of points in the $i$-th object. An extra one is added to avoid the denominator being zero. $F^{i}_{avg}$ represents the expected mean value of all the pixels in the $i$-th object, thereby we can compute the object loss $L_{obj}$ for all objects as:
@@
where $\mathcal{MSE}$ is the mean squared error function.
Clearly, the proposed $L_{obj}$ directly leverages the semantically meaningless regions generated by SAM, fully utilizing the detailed segmentation mask information in SGO.
III-B2 Boundary Loss
Previous studies [41, 42, 33] have demonstrated that incorporating edge constraints can effectively enhance the performance of segmentation models in remote sensing tasks. Our observations indicate that SGO inherently contains highly detailed boundary information, as depicted in Fig. 1 (d). To leverage this boundary information, we set the boundary to 00 in $Y_{o}$ and generate SGB, denoted as $Y_{b}$. In this work, the boundary metric ($BF_{1}$) [33], capable of directly computing boundary loss from the segmentation output $P$ of the semantic model, is employed to evaluate the precision of boundary detection. The boundary loss $L_{bdy}$ is given by:
$L_{bdy}=1-BF_{1}.,$ (6)
@@ III-B2 Boundary Loss
Finally, the overall objective function employed in training the segmentation model is given by Eq. (2), which sums the semantic segmentation loss $L_{seg}$, the object loss $L_{obj}$, and the boundary loss $L_{bdy}$.
@@ IV-A1 ISPRS Vaihingen
The ISPRS Vaihingen dataset comprises $16$ true orthophotos with very high-resolution, averaging $2500\times 2000$ pixels. Each orthophoto encompasses three channels, namely Near-Infrared, Red, and Green (NIRRG) of $9$ cm ground sampling distance. This dataset includes five foreground classes, namely impervious surface, building, low vegetation, tree, car, and one background class (clutter). The $16$ orthophotos are divided into a training set of $12$ patches, and a test set of $4$ patches. The training set comprises orthophotos of index numbers $1,3,23,26,7,11,13,28,17,32,34,37$, while the test set $5,21,15,30$.
@@ IV-A2 LoveDA Urban
The LoveDA dataset contains two scenes, namely Urban and Rural. Considering the diverse distribution of ground objects, we selected the LoveDA Urban scene for our experiments. The LoveDA Urban comprises $1833$ high-resolution optical remote sensing images, each with the size of $1024\times 1024$ pixels. The images provide three channels, namely Red, Green, and Blue (RGB), with a ground sampling distance (GSD) of $30$ cm. The dataset encompasses seven landcover categories, including background, building, road, water, barren, forest, and agriculture [43]. These images were collected from three cities (Nanjing, Changzhou, and Wuhan) in China. The $1833$ images are divided into two parts, with $1156$ images for training and $677$ images for testing. Specifically, the training set contains images indexed from $1366$ to $2521$, while the test set spans images from $3514$ to $4190$.
The two datasets differ in sampling resolution, ground object categories, and label accuracy. Notably, the ISPRS Vaihingen exhibits higher sampling accuracy, a reduced number of categories, and more precise ground truth. In contrast, the LoveDA Urban presents the opposite characteristics. Conducting experiments on both datasets provides substantial evidence regarding the effectiveness of our framework. Throughout the training and testing processes, a sliding window is employed for dynamically assembling training batches, enabling the processing of large images without pre-splitting. The sliding window size is set at $256\times 256$, with a stride of $256$ during training and an adjusted stride of $32$ during testing. The use of a smaller stride in testing effectively mitigates border effects by averaging the prediction outcomes within overlapping regions [44, 45].
TABLE I: Experimental results on the ISPRS Vaihingen dataset. We present the OA of five foreground classes and three overall performance metrics. Accuracy of each category is presented in the F1/IoU form. Bold values are the best.
baseline ResNet-18 92.33/85.76 96.25/92.78 80.47/67.33 90.85/83.22 89.35/80.75 89.85 81.97
Ours ResNet-18 92.79/86.54 96.74/93.69 80.85/67.86 91.17/83.77 90.43/82.53 90.40 82.88
baseline SwinTrans-base 93.41/87.64 96.92/94.02 81.53/68.82 90.91/83.33 88.46/79.31 90.24 82.62
Ours SwinTrans-base 93.73/88.20 97.19/94.53 81.77/69.16 91.62/84.53 91.09/83.64 91.08 84.01
Figure 3: Qualitative performance comparisons on the ISPRS Vaihaigen with the size of $512\times 512$. (a) NIRRG images, (b) Ground truth, (c) baseline, (d) Ours. The first two rows show the results on ResNet-18 while the last two rows show the results on SwinTrans-base. Some purple boxes are added to highlight the differences.
@@ IV-C Implementation details
The experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU equipped with 24GB RAM. To generate SGO and SGB, we utilize the interface offered by Meta AI, which involves three pertinent parameters, namely “crop_nms_thresh”, “box_nms_thresh”, and “pred_iou_thresh” whose definitions can be found in Meta AI documentation [^1] Meta AI: https://github.com/facebookresearch/segment-anything/tree/main/segment_anything. In our experiments, these values were configured to $0.5$, $0.5$, and $0.96$, respectively. Note that a higher “pred_iou_thresh” value tends to generate more dependable raw output. Stochastic gradient descent (SGD) was employed as the optimization algorithm for training all models. Furthermore, the experiments employed a learning rate of $0.01$, a momentum of $0.9$, a decaying coefficient of $0.0005$, and a batch size of $10$. The values of $\lambda_{o}$ and $\lambda_{b}$ are set to $1.0$ and $0.1$, respectively.
Simple data augmentations, such as random rotation and flipping, were utilized. In addition, UNetformer [14], which benefits from its seamless integration of both CNN and transformer encoders, was employed as our baseline. This choice allows us to validate our approach across different classical networks. Specifically, the former utilizes ResNet (ResNet-18) [6] as its backbone, while the latter employs the Swin Transformer (SwinTrans-base) [46].
TABLE II: Experimental results on the LoveDA Urban dataset. We present the OA of five foreground classes and three overall performance metrics. Accuracy of each category is presented in the F1/IoU form. Bold values are the best.
@@
baseline ResNet-18 54.66/37.61 69.09/52.78 68.33/51.89 77.66/63.47 56.98/39.84 51.01/34.23 20.54/11.44 65.34 49.12
Ours ResNet-18 55.43/38.34 75.91/61.17 73.61/58.24 80.55/67.43 43.19/27.55 57.05/39.91 43.64/27.91 65.74 50.55
baseline SwinTrans-base 55.38/38.30 75.46/61.21 75.29/60.38 75.81/61.04 40.83/25.65 54.10/37.08 40.32/25.25 64.95 49.71
Ours SwinTrans-base 56.40/39.27 77.23/62.91 74.74/59.67 77.30/63.00 44.44/28.57 53.30/36.33 44.17/28.34 66.02 50.68
Figure 4: Qualitative performance comparisons on the LoveDA Urban with the size of $512\times 512$. (a) RGB images, (b) Ground truth, (c) baseline, (d) Ours. The first two rows show the results on ResNet-18 while the last two rows show the results on SwinTrans-base. Some purple boxes are added to highlight the differences.
@@ IV-D Performance Comparison
We benchmarked the performance of the proposed framework on two representative backbones, namely ResNet-18 and SwinTrans-base. The quantitative results are listed in Table I and Table II.
@@ IV-D1 Performance Comparison on the Vaihingen dataset
As indicated in Table I, our proposed framework demonstrates significant improvements in terms of both mF1 and mIoU metrics as compared to the baseline method. These results affirm the effectiveness of our approach in robustly capturing object and boundary representations. Notably, the incorporation of SGO and SGB information substantially enhances semantic segmentation performance in remote sensing images. When compared to the baseline, our method outperformed across all five classes, namely impervious surface, building, low vegetation, tree, and car. In particular, on the ISPRS Vaihingen dataset, ours (ResNet-18) achieved performance improvement of $1.08\%$ and $1.78\%$ in F1 and IoU, respectively, on the Car class as compared to baseline. Furthermore, the classification accuracy for building was enhanced by $0.49\%$ and $0.91\%$ in F1 and IoU, respectively, compared to the baseline. These improvements can be attributed to the inherent characteristics of car and building classes, which typically exhibit simpler and more standardized shapes compared to other ground object categories. For example, a car often presents a fixed-size rectangular shape under consistent sampling conditions, while a building typically manifests as a changing-size rectangle. As a result, SGO and SGB provide more reliable criteria for these specific categories, facilitating more accurate segmentation.
@@
In contrast, impervious surface, low vegetation, and tree often feature intricate boundaries and varying sizes. Despite their complexity, SGO and SGB still offer valuable supplementary information for these specific categories. For instance, the classification accuracy for impervious surface was improved by $0.46\%$ and $0.78\%$ in F1 and IoU, respectively. Meanwhile, both low vegetation and tree showed comparable improvements in F1 and IoU metrics, reflecting their similar characteristics. In terms of overall performance, our model achieved an mF1 score of $90.40\%$ and an mIoU of $82.88\%$, representing an increase of $0.55\%$ and $0.91\%$ compared to the baseline, respectively.
@@
Our experiments with SwinTrans-base further substantiated our observations above. Specifically, our model (SwinTrans-base) demonstrated significant enhancements, exhibiting a notable $2.63\%$ and $4.33\%$ improvement in F1 and IoU metrics, respectively, for the Car class compared to the baseline. Moreover, we observed a slight increase in classification accuracy for the building, with improvements of $0.27\%$ and $0.51\%$ in F1 and IoU, respectively. In particular, the tree category exhibits a notable improvement, while the improvement in the low vegetation category is somewhat diminished, primarily due to their significant similarity. The overall performance corresponds to an mF1 score of $91.08\%$ and an mIoU of $84.01\%$, marking respective increases of $0.84\%$ and $1.39\%$ compared to the baseline. These outcomes affirm that our proposed framework, with the aid of SGO and SGB across various models, achieved superior generalization performance.
@@
TABLE III: Ablation studies on the ISPRS Vaihingen dataset. Accuracy of each category is presented in the F1/IoU form. Bold values are the best.
@@
TABLE IV: Ablation studies on the LoveDA Urban dataset. Accuracy of each category is presented in the F1/IoU form. Bold values are the best.
@@
Fig. 3 presents a visual comparison between the results produced by the baseline method and our approach. Throughout all subfigures, purple boxes highlight areas of interest. Firstly, our method excelled in precisely segmenting entire objects, notably evident in the clear delineation of building shown in the first and fourth rows. Secondly, our approach significantly refined object boundaries, as seen low vegetation in the second row and the building in the third row, aligning more closely with the actual ground truth boundaries. Furthermore, our method successfully addressed the segmentation of challenging categories, such as tree in the second row, building in the third row, and impervious surface and car in the fourth row. These improvements primarily stem from the object region and boundary details contained in the SGO and SGB. Leveraging the well-designed framework founded on object and boundary loss functions allows for the comprehensive exploitation of this detailed information, leading to the observed improvements in segmentation accuracy.
IV-D2 Performance Comparison on the LoveDA Urban
Experiments conducted on the LoveDA Urban dataset yielded results similar to those observed in the ISPRS Vaihingen dataset, despite variations in sampling resolution and ground object categories between the two datasets. As shown in Table II, the classification accuracy for building, water, and agriculture were $75.91\%/61.17\%$, $80.55\%/67.43\%$, and $43.64\%/27.91\%$, respectively, which amounts to an increase of $6.82\%/8.39\%$, $2.89\%/3.96\%$, and $23.1\%/16.47\%$ in F1 and IoU, respectively, as compared to the baseline (ResNet-18). Meanwhile, our findings with SwinTrans-base consistently confirmed advantages for these classes characterized by regular shapes and uncomplicated borders. The corresponding mF1 and mIoU values were $65.74\%$, $50.55\%$ respectively, marking increases of $0.4\%$, and $1.43\%$, respectively, over baseline (ResNet-18) while the increases for mF1 and mIoU on SwinTrans-base were $1.07\%$, $0.97\%$.
However, upon evaluating this dataset, we observed fluctuating performance changes when comparing the results of the two networks. ResNet-18 exhibited improved performance for road and forest, whereas SwinTrans-base demonstrated the opposite trend. Conversely, the performance for barren decreased in ResNet-18 but improved in SwinTrans-base. The primary reason for this discrepancy is that these three categories have intricate boundaries and varying sizes. Notably, the proportion of barren is very small, leading to volatility in its results during the experiments [43]. While the segmentation model aims to enhance overall performance, it cannot guarantee consistent improvement in every category. Nevertheless, the overall performance improvement of our framework demonstrates the instructive nature of this approach for subsequent works.
Fig. 4 presents a visualization example from the LoveDA Urban, with purple boxes highlighting areas of interest in all subfigures. Firstly, the model demonstrated enhanced accuracy in identifying building, evident across the first, second, and third rows. Furthermore, a more complete identification of water and agriculture objects was apparent, exemplified by the water in the first row and agriculture in the second and fourth rows. These observations closely align with the improvements indicated by the mF1 and mIoU metrics as presented in Table II.
IV-E Ablation Study

### 20 Dec 2023 v2

@@ III-B1 Object Consistency Loss
where $\mathcal{G}$ calculates the sum of all pixels in the spatial dimension and reshapes to its original shape while $N^{i}$ is the number of points in the $i$-th object. An extra one is added to avoid the denominator being zero. $F^{i}_{avg}$ represents the expected mean value of all the pixels in the $i$-th object. Thus, we can compute the object consistency loss $L_{obj}$ for all objects as:
@@
where $\mathcal{MSE}(\cdot)$ is the mean squared error function.
Clearly, the proposed $L_{obj}$ directly leverages the regions generated by SAM, fully utilizing the detailed segmentation mask information in SGO.
III-B2 Boundary Preservation Loss
Previous studies [61, 62, 42] have demonstrated that incorporating edge constraints can effectively enhance the performance of semantic segmentation models in remote sensing tasks. Our observations indicate that SGO inherently contains highly detailed boundary information, as depicted in Fig. 1 (d). To leverage this boundary information, we set the boundary to 00 in $Y_{o}$ and generate SGB, denoted as $Y_{b}$. In this work, the boundary metric ($BF_{1}$) [42], capable of directly computing boundary preservation loss from the segmentation output $P$ of the semantic model, is employed to evaluate the precision of boundary detection. The boundary preservation loss $L_{bdy}$ is given by:
$L_{bdy}=1-BF_{1},$ (6)
@@ III-B2 Boundary Preservation Loss
Finally, the overall objective function employed in training the semantic segmentation model is given by Eq. (2), which sums the semantic segmentation loss $L_{seg}$, the object consistency loss $L_{obj}$, and the boundary preservation loss $L_{bdy}$.
TABLE I: Experimental results on the ISPRS Vaihingen dataset. We present the OA of five foreground classes and three overall performance metrics. The accuracy of each category is presented in the F1/IoU form. Bold values are the best.
ABCNet [63] ResNet-18 89.78/81.45 94.30/89.21 78.49/64.59 90.08/81.95 74.05/58.80 85.34 75.20
ABCNet+SAM ResNet-18 92.02/85.21 95.97/92.25 79.90/66.53 90.06/81.92 78.52/64.64 87.29 78.11
CMTFNet [24] ResNet-50 92.53/86.09 96.95/94.09 79.98/66.64 90.22/82.19 89.87/81.60 89.91 82.12
CMTFNet+SAM ResNet-50 92.98/86.88 97.05/94.26 80.58/67.48 91.19/83.81 90.98/83.44 90.56 83.18
UNetformer [21] ResNet-18 92.33/85.76 96.25/92.78 80.47/67.33 90.85/83.22 89.35/80.75 89.85 81.97
UNetformer+SAM ResNet-18 92.79/86.54 96.74/93.69 80.85/67.86 91.17/83.77 90.43/82.53 90.40 82.88
FTUNetformer [21] SwinTrans-base 93.41/87.64 96.92/94.02 81.53/68.82 90.91/83.33 88.46/79.31 90.24 82.62
FTUNetformer+SAM SwinTrans-base 93.73/88.20 97.19/94.53 81.77/69.16 91.62/84.53 91.09/83.64 91.08 84.01
@@ IV-A1 ISPRS Vaihingen
The ISPRS Vaihingen dataset comprises $16$ true orthophotos with very high-resolution, averaging $2500\times 2000$ pixels. Each orthophoto encompasses three channels, namely Near-Infrared, Red, and Green (NIRRG) of $9$ cm ground sampling distance. This dataset includes five foreground classes, namely impervious surface, building, low vegetation, tree, car, and one background class (clutter). The $16$ orthophotos are divided into a training set of $12$ patches and a test set of $4$ patches. The training set comprises orthophotos of index numbers $1,3,23,26,7,11,13,28,17,32,34,37$, while the test set $5,21,15,30$.
@@ IV-A2 LoveDA Urban
The LoveDA dataset contains two scenes, namely Urban and Rural. Considering the diverse distribution of ground objects, we selected the LoveDA Urban scene for our experiments. The LoveDA Urban comprises $1833$ high-resolution optical remote sensing images, each with the size of $1024\times 1024$ pixels. The images provide three channels, namely Red, Green, and Blue (RGB), with a ground sampling distance of $30$ cm. The dataset encompasses seven landcover categories, including background, building, road, water, barren, forest, and agriculture [64]. These images were collected from three cities in China (Nanjing, Changzhou, and Wuhan). The $1833$ images are divided into two parts, with $1156$ images for training and $677$ images for testing. Specifically, the training set contains images indexed from $1366$ to $2521$, while the test set spans images from $3514$ to $4190$.
The two datasets differ in sampling resolution, ground object categories, and label accuracy. Notably, the ISPRS Vaihingen exhibits higher sampling accuracy, a reduced number of categories, and more precise ground truth. In contrast, the LoveDA Urban presents the opposite characteristics. Conducting experiments on both datasets provides substantial evidence regarding the effectiveness of our framework. Throughout the training and testing processes, a sliding window is employed for dynamically assembling training batches, enabling the processing of large images without pre-splitting. The sliding window size is set to $256\times 256$, with a stride of $256$ during training and an adjusted stride of $32$ during testing. The use of a smaller stride in testing effectively mitigates border effects by averaging the prediction outcomes within overlapping regions [65, 66].
Figure 6: Qualitative performance comparisons on the ISPRS Vaihaigen with the size of $512\times 512$. (a) ABCNet, (b) CMTFNet, (c) UNetformer, (d) FTUNetformer. We showcase two samples for each model. Some purple boxes are added to highlight the differences.
@@ IV-C Implementation details
The experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU equipped with 24GB RAM. To generate SGO and SGB, we utilize the interface offered by Meta AI, which involves three pertinent parameters, namely “crop_nms_thresh”, “box_nms_thresh”, and “pred_iou_thresh” whose definitions can be found in Meta AI documentation [^1] Meta AI: https://github.com/facebookresearch/segment-anything/tree/main/segment_anything. In our experiments, these values were configured to $0.5$, $0.5$, and $0.96$, respectively. Note that a higher “pred_iou_thresh” value tends to generate more dependable raw output. The threshold $K$ and $S$ are both set to $50$. Stochastic gradient descent (SGD) was employed as the optimization algorithm for training all models. Furthermore, the experiments employed a learning rate of $0.01$, a momentum of $0.9$, a decaying coefficient of $0.0005$, and a batch size of $10$. The values of $\lambda_{o}$ and $\lambda_{b}$ were chosen based on our sensitivity analysis as shown in Section IV-E. Simple data augmentations, such as random rotation and flipping, were utilized for all experiments.
TABLE II: Experimental results on the LoveDA Urban dataset. We present the OA of five foreground classes and three overall performance metrics. The accuracy of each category is presented in the F1/IoU form. Bold values are the best.
@@
ABCNet [63] ResNet-18 52.02/35.15 63.36/46.37 65.42/48.61 61.42/44.31 44.27/28.43 54.63/37.58 19.98/11.10 57.30 40.58
ABCNet+SAM ResNet-18 50.23/33.54 68.25/51.80 66.10/49.36 75.74/60.95 36.09/22.01 55.43/38.34 22.59/12.74 59.28 43.53
CMTFNet [24] ResNet-50 56.09/38.98 74.18/58.96 67.11/50.50 70.3/554.27 47.00/30.72 54.45/37.41 41.92/25.65 62.95 46.68
CMTFNet+SAM ResNet-50 55.86/38.76 75.79/61.02 76.05/61.36 72.14/56.42 37.28/22.91 55.97/38.86 43.24/27.58 63.43 48.09
UNetformer [21] ResNet-18 54.66/37.61 69.09/52.78 68.33/51.89 77.66/63.47 56.98/39.84 51.01/34.23 20.54/11.44 65.34 49.12
UNetformer+SAM ResNet-18 55.43/38.34 75.91/61.17 73.61/58.24 80.55/67.43 43.19/27.55 57.05/39.91 43.64/27.91 65.74 50.55
FTUNetformer [21] SwinTrans-base 55.38/38.30 75.46/61.21 75.29/60.38 75.81/61.04 40.83/25.65 54.10/37.08 40.32/25.25 64.95 49.71
FTUNetformer+SAM SwinTrans-base 56.40/39.27 77.23/62.91 74.74/59.67 77.30/63.00 44.44/28.57 53.30/36.33 44.17/28.34 66.02 50.68
@@ IV-D Performance Comparison
We benchmarked the performance of the proposed framework on four representative semantic segmentation models for remote sensing images, namely ABCNet [63], CMTFNet [24], UNetformer [21] and FTUNetformer [21]. Specifically, these four methods employ the most common backbones recently, the CNN-based ResNet [10] and the self-attention-based [11] Swin Transformer [67]. This choice allows us to validate our approach across different classical networks. The quantitative results are listed in Table I and Table II.
@@ IV-D1 Performance Comparison on the Vaihingen dataset
As indicated in Table I, our proposed framework demonstrates significant improvements in terms of both mF1 and mIoU metrics as compared to the four baseline methods. These results affirm the effectiveness of our approach in robustly capturing object and boundary representations. Notably, the incorporation of SGO and SGB information substantially enhances semantic segmentation performance in remote sensing images. For illustration purposes, we focus on the third set of experiments in which UNetformer+SAM demonstrated superior performance over UNetformer in all five classes, namely impervious surface, building, low vegetation, tree, and car. In particular, on the ISPRS Vaihingen dataset, UNetformer+SAM achieved performance improvement of $1.08\%$ and $1.78\%$ in F1 and IoU, respectively, on the Car class as compared to UNetformer. Furthermore, the classification accuracy for building was enhanced by $0.49\%$ and $0.91\%$ in F1 and IoU, respectively, compared to UNetformer. These improvements can be attributed to the inherent characteristics of car and building classes, which typically exhibit simpler and more standardized shapes compared to other ground object categories. For instance, a car is often presented as a fixed-size rectangular shape under consistent sampling conditions, while a building typically manifests as a changing-size rectangle. As a result, SGO and SGB provide more reliable criteria for these specific categories, facilitating more accurate segmentation.
@@
In contrast, impervious surface, low vegetation, and tree often feature intricate boundaries and varying sizes. Despite their complexity, SGO and SGB still offer valuable supplementary information for these specific categories. For instance, the classification accuracy for impervious surface was improved by $0.46\%$ and $0.78\%$ in F1 and IoU, respectively. Meanwhile, both low vegetation and tree showed comparable improvements in F1 and IoU metrics, reflecting their similar characteristics. In terms of overall performance, UNetformer+SAM achieved an mF1 score of $90.40\%$ and a mIoU of $82.88\%$, representing an increase of $0.55\%$ and $0.91\%$ compared to UNetformer, respectively.
@@
Other comparative experiments in Table I further substantiated our observations above. Specifically, FTUNetformer+SAM that adopts Swin Transformer as its backbone, also demonstrated significant enhancements, exhibiting a notable $2.63\%$ and $4.33\%$ improvement in F1 and IoU metrics, respectively, for the Car class compared to FTUNetformer. Moreover, we observed a slight increase in classification accuracy for the building, with improvements of $0.27\%$ and $0.51\%$ in F1 and IoU, respectively. In particular, the tree category exhibits a notable improvement, while the improvement in the low vegetation category is somewhat diminished, primarily due to their significant similarity. The overall performance corresponds to an mF1 score of $91.08\%$ and an mIoU of $84.01\%$, marking respective increases of $0.84\%$ and $1.39\%$ compared to the FTUNetformer. These outcomes affirm that our proposed framework, with the aid of SGO and SGB, achieved superior generalization performance across various models.
Figure 7: Qualitative performance comparisons on the LoveDA Urban with the size of $512\times 512$. (a) ABCNet, (b) CMTFNet, (c) UNetformer, (d) FTUNetformer. We showcase two samples for each model. Some purple boxes are added to highlight the differences.
Fig. 6 presents a visual comparison between the results produced by the baseline methods and our approach. Throughout all subfigures, purple boxes highlight areas of interest. Firstly, our method excelled in precisely segmenting entire objects, notably evident in the clear delineation of building. Secondly, our approach significantly refined object boundaries, as seen low vegetation and building, aligning more closely with the actual ground truth boundaries. Furthermore, our method successfully addressed the segmentation of challenging categories, such as tree, building and impervious surface. These improvements primarily stem from the object region and boundary details contained in the SGO and SGB. Leveraging the well-designed framework founded on object- and boundary-based loss functions allows for the comprehensive exploitation of these detailed information, leading to the observed improvements in segmentation accuracy.
IV-D2 Performance Comparison on the LoveDA Urban
Experiments conducted on the LoveDA Urban dataset yielded results similar to those observed in the ISPRS Vaihingen dataset despite variations in sampling resolution and ground object categories between the two datasets. Given the modest benchmark performance of ABCNet and CMTFNet, our approach yielded notably superior enhancements in overall performance. Observing the results of UNetformer and UNetformer+SAM as shown in Table II, the classification accuracy for building, water, and agriculture were $75.91\%/61.17\%$, $80.55\%/67.43\%$, and $43.64\%/27.91\%$, respectively, which amounts to an increase of $6.82\%/8.39\%$, $2.89\%/3.96\%$, and $23.1\%/16.47\%$ in F1 and IoU, respectively, as compared to UNetformer. Meanwhile, our findings with FTUNetformer+SAM consistently confirmed advantages for these classes characterized by regular shapes and uncomplicated borders. The corresponding mF1 and mIoU values were $65.74\%$, $50.55\%$ respectively, marking increases of $0.4\%$, and $1.43\%$, respectively, over UNetformer while the increases for mF1 and mIoU on FTUNetformer were $1.07\%$, $0.97\%$.
However, upon evaluating this dataset, we observed fluctuating performance changes when comparing the results of the four baselines. For instance, UNetformer+SAM exhibited improved performance for road and forest, whereas FTUNetformer+SAM demonstrated the opposite trend. Conversely, the performance for barren decreased in UNetformer+SAM but improved in FTUNetformer+SAM. The primary reason for this discrepancy is that these three categories have intricate boundaries and varying sizes. Notably, the proportion of barren is very small, leading to volatility in its results during the experiments [64]. While the semantic segmentation model aims to enhance overall performance, it cannot guarantee consistent improvement in every single category. Nevertheless, the overall performance improvement of our framework demonstrates the instructive nature of this approach for subsequent remote sensing works with different models.
Fig. 8 presents visualization examples from the LoveDA Urban, with purple boxes highlighting areas of interest in all subfigures. Firstly, the model demonstrated enhanced accuracy in identifying building, evident across all methods. Furthermore, a more complete identification of water and agriculture objects was apparent. These observations closely align with the improvements indicated by the mF1 and mIoU metrics as presented in Table II.
Figure 8: Sensitivity analysis of (a) $\lambda_{o}$ and (b) $\lambda_{b}$ on the ISPRS Vaihingen dataset. The performance of the baseline is depicted using dashed lines to highlight the variation of performance.
@@ IV-D2 Performance Comparison on the LoveDA Urban
TABLE III: Ablation studies on the ISPRS Vaihingen dataset. The accuracy of each category is presented in the F1/IoU form. Bold values are the best.
@@
TABLE IV: Ablation studies on the LoveDA Urban dataset. The accuracy of each category is presented in the F1/IoU form. Bold values are the best.
@@
TABLE V: Computational complexity analysis measured by two $256\times 256$ images on a single NVIDIA GeForce RTX 4090 GPU. MIoU values are the results on the ISPRS Vaihingen dataset. Bold values are the best.
Model FLOPs (G) Parameter (M) Memory (MB) Running Time ($10^{-3}$s) Inference Speed (FPS) MIoU(%)
ABCNet [63]
7.81 13.67 2480 1.23 32.62
75.20
ABCNet+SAM
7.81 13.67 2480
4.99
32.62 78.11
CMTFNet [24]
17.14 30.07 3240 1.84 14.47
82.12
CMTFNet+SAM
17.14 30.07 3240
5.97
14.47 83.18
UNetformer [21]
5.87 11.69 2490 14.35 31.11
81.97
UNetformer+SAM
5.87 11.69 2490
50.56
31.11 82.88
FTUNetformer [21]
50.84 96.14 5176 3.40 21.35
82.62
FTUNetformer+SAM
50.84 96.14 5176
74.85
21.35 84.01
IV-E Sensitivity Analysis
The analysis focuses on the sensitivity of two hyper-parameters, $\lambda_{o}$ and $\lambda_{b}$. $\lambda_{o}$ is designed to adjust the influence stemming from object consistency, whereas $\lambda_{b}$ adjusts the contribution originating from boundary information. These parameters hold significance in balancing the object consistency loss and boundary preservation loss. Considering that there is no strong correlation between these hyper-parameters, separate sensitivity experiments were conducted using UNetformer.
Fig. 8 (a) illustrates the model performance in terms of mF1 and mIoU across varying $\lambda_{o}$ values. A smaller $\lambda_{o}$ diminishes the influence of object information, while a larger value can overly emphasize its significance. Notably, setting $\lambda_{o}\geq 2.0$ resulted in a noticeable degradation in performance. Conversely, within the range $\lambda_{o}\in\left[0.1,1.0\right]$, performance exhibited lower sensitivity to changes in $\lambda_{o}$. Hereby our experiments employed $\lambda_{o}=1.0$, unless specified otherwise. Additionally, in Fig. 8 (b), the model’s performance is depicted concerning varying $\lambda_{b}\in\left[0.01,1.0\right]$. A smaller $\lambda_{b}$ value made the contribution from the boundary information insignificant in learning the model. The best performance was observed when $\lambda_{b}$ reached $0.1$, as illustrated in Fig. 8 (b). However, further increments in $\lambda_{b}$ resulted in performance degradation. Ultimately, our experiments set $\lambda_{o}$ and $\lambda_{b}$ to $1.0$ and $0.1$, respectively, aligning them within a similar order of magnitude.
IV-F Ablation Study